OpenStudy (anonymous):

e^.05t=1600 how to solve?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[e^{.05t}=1600\] \[.05t=ln(1600)\]\ \[t=\frac{ln(1600)}{.05}\]

OpenStudy (anonymous):

so the answer is t=In(1600)/.05?

OpenStudy (anonymous):

the answer is t =14.75

OpenStudy (anonymous):

use caculator

OpenStudy (anonymous):

I did. & it's right. thanks :)

OpenStudy (anonymous):

u r w c

OpenStudy (anonymous):

how did you get 14.75?

OpenStudy (anonymous):

i got 147.55

OpenStudy (anonymous):

and so it is!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

In(4x)=3 c anyone help?

OpenStudy (anonymous):

\[\ln (4x)=3\]

OpenStudy (anonymous):

rewrite in exponential form \[4x=e^3\] \[x=\frac{e^3}{4}\]

OpenStudy (anonymous):

i got .462

OpenStudy (anonymous):

was it \[ln(4x)=3\]?

OpenStudy (anonymous):

ok so x= e^3/4?

OpenStudy (anonymous):

6.695 rounded

OpenStudy (anonymous):

whch would be x= 2.7183^3

OpenStudy (anonymous):

yes \[ln(x)=y\] same as \[x=e^y\]

OpenStudy (anonymous):

so \[ln(4x)=3\] same as \[4x=e^3\]

OpenStudy (anonymous):

i keep getting 5.0225

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

how do you solve 4+5e^-x=0